Two resistances, one 12 Ω and the other 18 Ω, are connected in parallel. What is the equivalent resistance of the parallel combination?


1
Expert's answer
2022-04-11T12:02:32-0400

The equivalent resistance of the parallel combination

"R=\frac{R_1R_2}{R_1+R_2}"

"R=\frac{12*18}{12+18}=7.2\:\Omega"